OpenStudy (anonymous):

Consider separate 1.0-L samples of He(g) and UF6(g), both at 1.00 atm and containing the same number of moles. What ratio of temperatures for the two samples would produce the same root mean square velocity? Consider separate 1.0-L samples of He(g) and UF6(g), both at 1.00 atm and containing the same number of moles. What ratio of temperatures for the two samples would produce the same root mean square velocity? @Chemistry

OpenStudy (anonymous):

\[V = \sqrt{3*R*T/MW}\] V= RMS velocity R=8.3145 J/K*mole T= Temperature K MW = Molecular weight in Kg

OpenStudy (anonymous):

use this information